# SeFFeC: Semantic Facial Feature Control for Fine-grained Face Editing

### v1

@@ 5 Conclusion
We proposed SeFFeC – a novel method that enables users to perform fine-grained and deterministic face shape editing with little effort. The key idea behind SeFFeC is the use of 23 semantic face features, which are determined from different groups of facial landmarks. Unlike existing methods, landmarks have the added advantage that they allow to calculate the values of these face features dynamically, which allows training SeFFeC without requiring any manual labelling. We also introduced a novel semantic face feature loss that encourages the model to manipulate only the desired face features while keeping unrelated features unchanged. Given that these features are continuous, we trained a scaling network that could learn how to scale the manipulation vector to achieve the desired change in facial appearance. This represents a significant advance given that, using prior methods, users had to move towards the desired appearance via trial and error. In contrast, SeFFeC allows for deterministic face editing, which paves the way for future algorithmic manipulations using SeFFeC as part of automated face animation pipelines.

### v3

@@ 5 Conclusion
We proposed UP-FacE – a novel method that enables user-predictable and precise face shape editing with little effort. The key idea behind UP-FacE is the use of 23 semantic face features, determined from different facial landmarks groups. Unlike existing methods, landmarks have the advantage that they allow the calculation of the values of these face features dynamically, which allows training UP-FacE without requiring any manual labelling. We also introduced a novel semantic face feature loss that encourages the model to manipulate only the desired face features while keeping unrelated features unchanged. Given that these features are continuous, we trained a scaling network that could learn how to scale the manipulation vector to achieve the desired change in facial appearance. This represents a significant advance given that with prior methods, users had to move towards the desired appearance via trial and error. In contrast, UP-FacE allows for user-predictable face editing, which paves the way for more user-friendly digital face editing tools.